Find the resistance of a 12 meter length of metal wire of 0.5 mm diameter. The resistivity of metal is 2.8×10-6 Ω·cm at a temperature of 20°C. Express the answer with two decimal places.
Answer #1

L = length of wire = 12 m

d = diameter = 0.5 mm = 0.5 x 10-3 m

A = area of cross-section = \pid2/4 = (3.14) (0.5 x 10-3 )2 /4 = 1.9625 x 10-7 m2

\rho = resistivity = 2.8 x 10-6 ohm-cm = 2.8 x 10-8 ohm-m

resistance is given as

R = \rho L/A = (2.8 x 10-8 ) (12) / (1.9625 x 10-7 ) = 1.71 ohm